# A Flat Minima Perspective on Understanding Augmentations and Model Robustness

### v1

@@ C.1.3 Results on Theorem 1’s Bound
C.1.4 Flatness Evaluations on (tiny)ImageNet
Table 4: Flatness metrics of ERM and different augmentations on CIFAR benchmarks (:absent:$\downarrow$:↓ : The lower the better, i.e., flatter minimum). The results are consistent with Table 1; All the augmentations populating meaningful representations near the original found flatter minima compared to ERM (AugMix, RandAug, PixMix, DeepAug.)
Dataset Metrics      ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug tinyImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 137.97 113.97 (-24.00) 124.82 (-13.15) 113.97 (-24.00) 137.97 (-0.00) 97.09 (-16.88) LPF $\downarrow$ 1.91 1.10 (-0.81) 1.10 (-0.81) 1.05 (-0.86) 1.53 (-0.38) 1.47 (-0.44) $\epsilon_{\text{sharp}}$downarrow 1.46 0.64 (-0.78) 0.77 (-0.69) 0.55 (-0.91) 0.56 (-0.90) 1.95 (+0.49) ImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 481.93 481.93 (-0.00) 470.81 (-11.12) 405.52 (-76.41) 405.52 (-76.41) 694.44 (+212.51) LPF $\downarrow$ 0.36 0.35 (-0.01) 0.40 (+0.04) 0.35 (-0.01) 0.44 (+0.08) 3.39 (+3.03) ϵsharpsubscriptitalic-$\epsilon_{\text{sharp}}\downarrow$ 47.70 45.68 (-2.02) 46.82 (-0.88) 40.76 (-6.94) 44.13 (-3.57) 211.67 (+163.97)
Training Details. We train the ResNet18 model as described in Section C.1.2. We train using stochastic gradient descent (SGD) initialized with a 0.1 learning rate, 0.0005 weight decay, and 0.9 momentum. The learning rate then follows a cosine decay schedule, beginning at 0.1 with the same 0.0005 weight-decay setting. We have deployed the same configuration for measuring flatness as described in Section C.1.1 except for tinyImageNet experiment deploying $\tau=0.5,\epsilon=10.0$ for numerical stability.
C.1.5 Clean Accuracies of Models (Table 3)
@@
The tendency of the ’good’ and the ’bad’ augmentations remains the same as in Table 3, which aligns with our theorems, considering that generalization bound also applies to train-test data discrepancies (target data $\mathcal{T}$ being the test dataset.) On the other hand, there is a weak correlation between robustness and the clean accuracy. For instance, all methods have inferior clean accuracy compared to ERM in ImageNet, irrespective of their robustness. There are some cases where the clean accuracy increases but the robustness drops (e.g., StyleAug in CIFAR10), and some cases have dropped clean accuracy but improved robustness (e.g., PixMix in ImageNet). Also, some have severely dropped clean accuracy and robustness (e.g., StyleAug in CIFAR-100). We observe all four cases (better clean accuracy, better robustness / better clean accuracy, worse robustness / worse clean accuracy, better robustness / worse clean acc, worse robustness) exist.
C.1.6 Additional Experiments on Varying Backbone Networks
To validate our theoretical findings under differing backbone networks, we present additional experimental results on CIFAR-100 robustness benchmarks, extending results from Table 3. In CIFAR-100 benchmarks, AugMix, RandAug, and PixMix had nontrivial augmented sampled within the proximal region, leading to flatter minima and overall enhanced robustness. Contrarily, DeepAug and StyleAug had no representations near the original, failing to promote flatter minima and improved robustness. The tendency holds exactly the same for varying backbone networks as well. In below (Table  5), we demonstrate that our findings are consistent across different backbone networks [47, 23, 52].
@@ C.1.6 Additional Experiments on Varying Backbone Networks
Backbone Benchmark ERM      AugMix      RandAug      PixMix      DeepAug      StyleAug AllConvNet CIFAR-100-C $\downarrow$ 56.88 43.13 (-13.75) 47.52 (-9.36) 37.70 (-19.18) 62.00 (+5.12) 56.90 (+0.02) CIFAR-100-$\overline{\text{C}}$downarrow 20.33 9.84 (-10.49) 13.86 (-6.47) 14.29 (-6.04) 64.24 (+43.91) 38.64 (+18.31) CIFAR-100, $L_{2}$downarrow 93.96 86.51 (-7.45) 88.96 (-5.00) 74.93 (-19.03) 95.29 (+1.33) 96.56 (+2.60) CIFAR-100, $L_{\infty}$downarrow 99.95 99.90 (-0.05) 99.94 (-0.01) 99.64 (-0.31) 99.99 (+0.04) 99.71 (-0.24) DenseNet CIFAR-100-C $\downarrow$ 61.21 43.32 (-17.89) 48.54 (-12.67) 37.58 (-23.63) 61.47 (+0.26) 68.97 (+7.76) CIFAR-100-$\overline{\text{C}}$downarrow 23.10 8.87 (-14.23) 13.68 (-9.42) 14.27 (-8.83) 66.20 (+43.10) 43.90 (+20.80) CIFAR-100, $L_{2}$downarrow 99.91 97.88 (-2.03) 99.44 (-0.47) 93.19 (-6.72) 98.70 (-1.21) 99.76 (-0.15) CIFAR-100, $L_{\infty}$downarrow 100.00 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) 99.99 (-0.01) ResNext29 CIFAR-100-C $\downarrow$ 58.05 39.60 (-18.45) 43.49 (-14.56) 31.35 (-26.70) 61.50 (+3.45) 70.11 (+12.06) CIFAR-100-$\overline{\text{C}}$downarrow 13.09 3.85 (-9.24) 6.99 (-6.10) 2.64 (-10.45) 66.61 (+53.52) 22.78 (+9.69) CIFAR-100, $L_{2}$downarrow 98.72 91.68 (-7.04) 95.95 (-2.77) 91.64 (-7.08) 98.36 (-0.36) 99.67 (+0.95) CIFAR-100, $L_{\infty}$downarrow 100.00 99.94 (-0.06) 99.97 (-0.03) 99.90 (-0.10) 99.96 (-0.04) 99.98 (-0.02)
Appendix D Detailed Review of Related Works
D.1 Data Augmentations Strategies
Existing augmentation methods can be grouped into model-free, model-based, and policy-based algorithms, borrowing the taxonomy suggested by [53].
Model-free augmentations can be divided into single-image and multiple-image methods. Single-image augmentations not only include elementary image manipulations such as translation, rotation, and color jittering but also includes masking methods such as CutOut [11] and Hide-and-Seek [31]. Multi-image augmentations exploit two or more images to construct augmented images, including the mixing of the two disjoint images and labels, such as Mixup [59] and CutMix [57]. Recently, it is empirically observed that a few methods, AugMix [21] and PixMix [22], exhibit their superiority over others in terms of model robustness.
Model-based augmentations utilize pretrained models to generate augmented data. A number of methods that exploit generative models including CGAN [40], and their variants [12, 38, 2, 54] aim to aid data imbalance problems. Other augmentations, including DeepAugment [19], ANT [44], and StyleAug [24], directly aim to enhance the classifier’s robustness against common corruption, adversarial attacks, or domain shifts.
Policy-based augmentations focus on designing an automatic way to determine the optimal augmentation strategies by employing reinforcement learning or adversarial training. From a pioneering method AutoAugment [9] that utilizes reinforcement learning for finding the best augmentation strategies, subsequent works such as Fast AA [36], Faster AA [18], and RandAugment [10] aim to enhance both the efficiency of policy search and the model performance. Adversarial training-based augmentation strategies, including AdaTransform [49], Adversarial AA [60], and AugMax [51] leverage adversarial perturbations which maximally disturbs samples to be misclassified into other labels, finally leading to improve model robustness against unseen domains.
D.2 Augmentations and Model Robustness
[37] have provided a game-theoretic perspective on augmentations and introduces a new proxy metric for measuring common corruption robustness. [63] and [50] have theoretically found out that adversarial perturbations in the latent space can simulate worst-case distributional shifts in the data. [43] have empirically found out that Mixup and CutOut with model weight averaging is shown to improve adversarial robustness. [55] has shown that Mixup enlarges boundary thickness, the marginal space between differently labeled samples, which is deemed to be highly correlated with adversarial robustness and common corruptions. [56] interpret the augmentation-originated gains in model robustness by explaining that augmentations make deep models utilize both high and low frequency information of images so as to enhance model robustness against data corruptions. [41] and [1] have theoretically shown that utilizing unlabeled data in training can improve adversarial robustness. [19] have empirically found out that exploiting diverse augmentations together improves robustness against both adversarial and common corruptions.
Nonetheless, most of the prior interpretations of how augmentation contributes to model robustness are either confined to specific types of augmentations and robustness, or empirical analysis. None of the prior works generally explain how augmentations theoretically can improve model robustness across diverse distributional shifts.
D.3 Other Related Works
[61] theoretically show the duality between SAM and AT, and empirically found the conditions where SAM outperforms AT. Nonetheless, the study linked input–parameter duality only in binary linear classifiers with Gaussian data and basic adversarial-training (AT). On the other hand, we show that the duality survives in deep, multi-class models, arbitrary data distributions, and general augmentation family whose distribution is dense around the data. The findings of [61] cannot be extended to augmentation-driven robustness.
[27] investigates how to preserve fairness under distribution shifts by reinterpreting such shifts as equivalent perturbations to both model weights and inputs, employing a transportation function. However, their work focuses exclusively on group fairness with respect to demographic parity and equal opportunity, and only demonstrates the feasibility of this equivalence without further extensions.
[64] shows in binary classification that flatter minima mitigate the boundary tilting problem [48], and empirically finds these minima correlate with greater robustness to common corruptions. However, the theory remains limited, as the boundary tilting problem is closely tied to adversarial robustness and is not explored beyond binary classification.
Appendix E Further Discussions
E.1 Practical Applications of Our Theorems
@@
Benchmark AugMix DeepAug StyleAug DeepAug + AugMix StyleAug + AugMix DeepAug + StyleAug CIFAR-10-C $\downarrow$ 15.24 31.92 30.50 15.04 (-16.88) 19.02 (-11.48) 24.84 (-5.66) CIFAR-10-$\overline{\text{C}}$downarrow 1.73 17.99 9.34 1.32 (-16.67) 1.55 (-7.79) 4.12 (-5.22) CIFAR-100-C $\downarrow$ 42.64 62.34 70.92 43.53 (-18.81) 49.13 (-21.79) 53.50 (-17.42) CIFAR-100-$\overline{\text{C}}$downarrow 5.01 39.54 30.77 3.67 (-35.87) 3.70 (-27.07) 7.60 (-23.17) CIFAR-10, $L_{2}$downarrow 70.68 91.07 91.13 69.30 (-21.77) 81.07 (-10.06) 87.30 (-3.83) CIFAR-10, $L_{\infty}$downarrow 99.57 100.00 99.97 94.15 (-4.22) 96.34 (-3.33) 99.80 (-0.17) CIFAR-100, $L_{2}$downarrow 92.78 98.37 99.67 94.15 (-4.22) 96.34 (-3.33) 97.53 (-2.14) CIFAR-100, $L_{\infty}$downarrow 99.67 99.99 99.98 99.89 (-0.10) 99.95 (-0.03) 99.96 (-0.02)
@@
Our study also provides explanations on the behavior of adversarial training (AT), and rooms for the improvement. While AT is known to strengthen robustness to adversarial attacks, models trained with AT still break down when the test distribution deviates largely from the training data, as in common corruption case [20, 39]. We conjecture that this fragility arises because AT typically discovers small $b$-flat minima: these minima are flat enough to absorb minor, norm-bounded perturbations, yet they fail to cope with large distribution shifts. Hence, pairing AT with data augmentation schemes that generate samples both close to and far from the original data manifold (such as AugMix) can offset AT’s weakness under severe shifts. Below (Table 88), we report results obtained by training with AT (FGSM, $\epsilon=8/255$) and AT in conjunction with AugMix.
Table 7: Flatness evaluations of AT and AT + AugMix on CIFAR-10/100.
Dataset Metrics AT AT + AugMix CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 158.15 113.78 (-44.37) LPF $\downarrow$ 1.25 0.58 (-0.67) $\epsilon_{\text{sharp}}$downarrow 16.36 10.29 (-6.07) CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 207.90 161.89 (-46.01) LPF $\downarrow$ 2.95 2.11 (-0.84) ϵsharpsubscriptitalic-$\epsilon_{\text{sharp}}\downarrow$ 20.81 10.34 (-10.47)
Robustness Benchmark AT AT + AugMix CIFAR-10-C $\downarrow$ 24.12 15.39 (-8.73) Common CIFAR-10-$\overline{\text{C}}$downarrow 3.25 1.30 (-1.95) Corruption CIFAR-100-C $\downarrow$ 52.50 44.03 (-8.47) CIFAR-100-$\overline{\text{C}}$downarrow 7.36 3.74 (-3.62) CIFAR-10, $L_{2}$downarrow 69.24 62.86 (-6.38) Adversarial CIFAR-10, $L_{\infty}$downarrow 94.55 98.40 (+3.85) Attacks CIFAR-100, $L_{2}$downarrow 89.28 88.60 (-0.68) CIFAR-100, $L_{\infty}$downarrow 98.42 99.75 (+1.33)
E.2 Comparison with Flat Minima Optimizers
We establish a formal connection between data augmentation and robustness through the notion of flat minima. Specifically, we prove the duality between the input space and parameter space manifolds (Theorem 1) and show that the nonzero probability proximal region induced by augmentation (Assumption 1) is mapped to an equivalent region in parameter space, promoting flatter minima (Theorem 2). Bridging these augmentation-induced flat minima to improved robustness (Theorem 3) leverages prior results indicating that flat minima tighten generalization bounds [7, 62] against distribution shifts. Empirically, several studies already confirm that optimizers explicitly seeking flat minima enhance robustness [15, 5, 28, 33, 61]. Nevertheless, our findings are not trivial extensions of the prior works on flat optimizers. The experiments reveal that not every augmentation leads to flatter minima or greater robustness. Our work provides a necessary condition for an augmentation to foster flatter minima (non-zero proximal density). This property correlates strongly with flat minima and robustness in our experiments. Together with the proven duality between input and parameter manifolds under general theoretical setups, we provide a distinct and novel contribution apart from existing research on flat minima optimizations.
E.3 Any Other Modalities Beyond Visual Data?
Our main experiments handle visual data, more specifically, image classification tasks. Also, it might be unexplored that the key claims of this work can be further applied to other modalities, such as language, audio, or even sensory signals from devices. Our experiments seem to be limited to vision cases, but we believe that it is an important step forward to understand theoretical importance of augmentations in machine learning. Furthermore, we want to emphasize that our main theoretical claims do not explicitly rely on the characteristics of visual data, so it can be further extended to the other form of modalities. In addition, extending to other complicated task beyond classification can be an appropriate future research direction. In particular, sequential training such as natural language processing or decision based on reinforcement learning probably show distinct and unique theoretical understanding beyond classification.

### v3

@@ C.5. Results on Theorem 1’s Bound
C.6. Clean Accuracies of Models (Table 35)
@@
There is a weak correlation between robustness and the clean accuracy. For instance, all methods have inferior clean accuracy compared to ERM in ImageNet, irrespective of their robustness. There are some cases where the clean accuracy increases but the robustness drops (e.g., StyleAug in CIFAR10), and some cases have dropped clean accuracy but improved robustness (e.g., PixMix in ImageNet). Also, some have severely dropped clean accuracy and robustness (e.g., StyleAug in CIFAR-100). We observe all four cases (better clean accuracy, better robustness / better clean accuracy, worse robustness / worse clean accuracy, better robustness / worse clean acc, worse robustness) exist.
C.7. Confidence Intervals for the Main Tables
@@
Dataset Metrics     ERM    AugMix    RandAug    PixMix    DeepAug    StyleAug CIFAR-10 $\mu_{\text{PAC-Bayes}}\downarrow$ 168.83 $\pm$ 13.50 117.16 $\pm$ 3.15 110.80 $\pm$ 0.55 102.87 $\pm$ 4.18 67.74 $\pm$ 3.27 122.80 $\pm$ 11.68 LPF $\downarrow$ 1.43 $\pm$ 0.29 0.54 $\pm$ 0.01 0.45 $\pm$ 0.03 0.37 $\pm$ 0.01 0.72 $\pm$ 0.10 1.13 $\pm$ 0.12 $\epsilon_{\text{sharp}}\downarrow$ 40.90 $\pm$ 3.90 25.16 $\pm$ 4.32 24.47 $\pm$ 2.81 26.57 $\pm$ 8.34 55.63 $\pm$ 34.78 130.64 $\pm$ 1.69 CIFAR-100 $\mu_{\text{PAC-Bayes}}\downarrow$ 181.10 $\pm$ 9.53 155.74 $\pm$ 0.01 149.79 $\pm$ 8.41 134.58 $\pm$ 1.64 72.24 $\pm$ 2.47 129.02 $\pm$ 13.25 LPF $\downarrow$ 2.37 $\pm$ 0.16 1.84 $\pm$ 0.01 1.69 $\pm$ 0.02 1.42 $\pm$ 0.03 2.37 $\pm$ 0.23 3.95 $\pm$ 0.14 $\epsilon_{\text{sharp}}\downarrow$ 42.39 $\pm$ 4.68 39.32 $\pm$ 2.93 37.76 $\pm$ 3.32 33.14 $\pm$ 4.76 84.48 $\pm$ 4.33 302.81 $\pm$ 5.14
   Benchmarks ERM    AugMix    RandAug    PixMix    DeepAug    StyleAug CIFAR-10-C $\downarrow$ 30.54 $\pm$ 0.45 15.24 $\pm$ 0.27 19.65 $\pm$ 0.32 10.60 $\pm$ 0.18 31.92 $\pm$ 3.90 30.50 $\pm$ 1.00 CIFAR-10-$\overline{\text{C}}$ $\downarrow$ 31.35 $\pm$ 0.59 20.28 $\pm$ 0.32 20.64 $\pm$ 0.14 14.60 $\pm$ 0.16 36.75 $\pm$ 2.43 36.57 $\pm$ 0.02 CIFAR-100-C $\downarrow$ 59.04 $\pm$ 0.45 42.64 $\pm$ 0.16 46.59 $\pm$ 0.74 35.20 $\pm$ 0.22 62.34 $\pm$ 3.33 70.91 $\pm$ 0.88 CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 62.43 $\pm$ 1.16 48.38 $\pm$ 0.27 48.32 $\pm$ 0.41 40.20 $\pm$ 0.18 67.91 $\pm$ 2.31 76.56 $\pm$ 0.57 CIFAR-10, $L_{2}$ $\downarrow$ 77.61 $\pm$ 3.86 70.76 $\pm$ 1.05 76.15 $\pm$ 1.65 65.81 $\pm$ 2.07 91.18 $\pm$ 1.80 91.06 $\pm$ 0.39 CIFAR-10, $L_{\infty}$ $\downarrow$ 98.49 $\pm$ 1.64 99.10 $\pm$ 0.16 99.86 $\pm$ 0.02 99.46 $\pm$ 0.26 100.00 $\pm$ 0.00 99.98 $\pm$ 0.03 CIFAR-100, $L_{2}$ $\downarrow$ 98.73 $\pm$ 0.62 92.76 $\pm$ 0.19 96.06 $\pm$ 1.85 90.69 $\pm$ 2.55 98.44 $\pm$ 0.50 99.69 $\pm$ 0.02 CIFAR-100, $L_{\infty}$ $\downarrow$ 99.94 $\pm$ 0.09 99.67 $\pm$ 0.09 99.94 $\pm$ 0.11 99.69 $\pm$ 0.44 99.99 $\pm$ 0.02 99.98 $\pm$ 0.02
   Benchmarks ERM    AugMix    DeepAug    PixMix    RandAug    StyleAug tinyImageNet-C $\downarrow$ 74.76 $\pm$ 0.47 63.48 $\pm$ 0.28 56.14 $\pm$ 0.12 61.51 $\pm$ 0.28 66.92 $\pm$ 0.21 79.42 $\pm$ 1.19 tinyImageNet, $L_{2}$ $\downarrow$ 57.77 $\pm$ 0.57 61.84 $\pm$ 1.13 52.86 $\pm$ 0.34 58.38 $\pm$ 1.41 63.19 $\pm$ 0.74 83.13 $\pm$ 0.63 tinyImageNet, $L_{\infty}$ $\downarrow$ 99.94 $\pm$ 0.02 99.93 $\pm$ 0.04 99.95 $\pm$ 0.03 99.98 $\pm$ 0.02 99.98 $\pm$ 0.01 100.0 $\pm$ 0.00
Dataset Metrics     ERM    AugMix    DeepAug    PixMix    RandAug    StyleAug tinyImageNet $\mu_{\text{PAC-Bayes}}\downarrow$ 136.83 $\pm$ 2.05 109.18 $\pm$ 0.97 128.41 $\pm$ 2.36 106.26 $\pm$ 1.97 109.36 $\pm$ 0.97 102.72 $\pm$ 1.45 LPF $\downarrow$ 5.95 $\pm$ 0.27 3.81 $\pm$ 0.17 4.96 $\pm$ 0.12 3.40 $\pm$ 0.09 4.04 $\pm$ 0.11 4.33 $\pm$ 0.20 $\epsilon_{\text{sharp}}\downarrow$ 15.35 $\pm$ 0.51 13.09 $\pm$ 0.32 25.02 $\pm$ 1.11 10.72 $\pm$ 0.66 21.30 $\pm$ 1.92 35.40 $\pm$ 3.75
Table 10: 95% confidence intervals for CIFAR and tinyImageNet experiments. (Table 2356.)
C.8. Licensing and Computational Resources
Licenses
AugMix and RandAugment are under Apache 2.0 License. PixMix, DeepAugment, and StyleAugment are under MIT License.
Computational Resources for Experiments
[…]
D. Further Discussions
Practical Applications of Our Theorems
@@ D. Further Discussions
Benchmark AugMix DeepAug StyleAug DeepAug + AugMix StyleAug + AugMix DeepAug + StyleAug CIFAR-10-C $\downarrow$ 15.24 31.92 30.50 15.04 (-16.88) 19.02 (-11.48) 24.84 (-5.66) CIFAR-10-$\overline{\text{C}}$ $\downarrow$ 1.73 17.99 9.34 1.32 (-16.67) 1.55 (-7.79) 4.12 (-5.22) CIFAR-100-C $\downarrow$ 42.64 62.34 70.92 43.53 (-18.81) 49.13 (-21.79) 53.50 (-17.42) CIFAR-100-$\overline{\text{C}}$ $\downarrow$ 5.01 39.54 30.77 3.67 (-35.87) 3.70 (-27.07) 7.60 (-23.17) CIFAR-10, $L_{2}$ $\downarrow$ 70.68 91.07 91.13 69.30 (-21.77) 81.07 (-10.06) 87.30 (-3.83) CIFAR-10, $L_{\infty}$ $\downarrow$ 99.57 100.00 99.97 94.15 (-4.22) 96.34 (-3.33) 99.80 (-0.17) CIFAR-100, $L_{2}$ $\downarrow$ 92.78 98.37 99.67 94.15 (-4.22) 96.34 (-3.33) 97.53 (-2.14) CIFAR-100, $L_{\infty}$ $\downarrow$ 99.67 99.99 99.98 99.89 (-0.10) 99.95 (-0.03) 99.96 (-0.02)
@@
Our study also provides explanations on the behavior of adversarial training (AT), and rooms for the improvement. While AT is known to strengthen robustness to adversarial attacks, models trained with AT still break down when the test distribution deviates largely from the training data, as in common corruption case (IN-C7:2019; IN-C-bar:2021). We conjecture that this fragility arises because AT typically discovers small $b$-flat minima: these minima are flat enough to absorb minor, norm-bounded perturbations, yet they fail to cope with large distribution shifts. Hence, pairing AT with data augmentation schemes that generate samples both close to and far from the original data manifold (such as AugMix) can offset AT’s weakness under severe shifts (Table 78.)
Beyond the Visual Domain
Our experiments focus on image classification, yet the PSA condition and accompanying theorems do not rely on properties specific to images. We therefore expect the framework to extend to other modalities such as language, audio, or sensor data, where augmentation is commonplace. Applying the analysis to sequential tasks such as NLP or reinforcement learning is a natural next step and may reveal modality-specific nuances; we leave these directions for future work.